OpenStudy (anonymous):

What is the value of the expression? 34 + 7 • 62

OpenStudy (anonymous):

@TwilightRain1

OpenStudy (anonymous):

6^2

OpenStudy (anonymous):

solve the mulitplication first

OpenStudy (anonymous):

62 times 7 plus 34

OpenStudy (anonymous):

multiplication*

OpenStudy (anonymous):

try doing it on paper

OpenStudy (anonymous):

1476?

OpenStudy (anonymous):

468

OpenStudy (anonymous):

Its actually 6^2

OpenStudy (anonymous):

468 is my answer

OpenStudy (anonymous):

oh okay so 6 times 6 plus 34

OpenStudy (anonymous):

no 70

OpenStudy (anonymous):

yeah, its 468

OpenStudy (anonymous):

try it on paper

OpenStudy (anonymous):

do 7x62 first

OpenStudy (anonymous):

she said it was 6^2 not 62

OpenStudy (anonymous):

then + 34

OpenStudy (anonymous):

6² = 36 36 * 7 = 252 252 + 34 = 286

OpenStudy (anonymous):

ok wat too many people but its 6^2

OpenStudy (anonymous):

468

OpenStudy (anonymous):

the answer is 70

OpenStudy (anonymous):

way*

OpenStudy (anonymous):

34 plus 36 is 70

OpenStudy (anonymous):

6² = 36 36 * 7 = 252 252 + 34 = 286 this is the correct answer.

OpenStudy (anonymous):

Umm..

OpenStudy (anonymous):

okay is the problem 34+ 7 times 62 or 34 + 7 times 6^2

OpenStudy (anonymous):

it is 34 + 7 x 6²